Find the measure of Angle 3.

measure of angle 1 = x
measure of angle 2 = x-6
1
Expert's answer
2018-06-18T08:52:08-0400
Consider a triangle. The total measure of three angles of a triangle is 180 degrees.
Thus, the measure of angle 3 = 180 - x - (x - 6) = 186 - 2x